Simplify.
x+2/x divided by x-2/3x
Thank you

Answer by sdmmadam@yahoo.com(530)

Simplify.
x+2/x divided by x-2/3x
=[(x+2)/x]divided by (x-2)/(3x)
=[(x+2)/x]multiplied by(3x)/(x-2)
=[(x+2)/x]X[(3x)/(x-2)]
= 3(x+2)/(x-2)
Note: when division symbol is replaced by multiplication symbol the fraction that comes after the division symbol should be reciprocated. that is after putting the multiplication symbol put in the place of the original fraction [1/fraction] that is here {1/[(x-2)/(3x)]} = (3x)/(x-2)}
In simple language change nr into dr and dr into numerator!)
Some examples:
1)p divided by (p/m)
= p multiplied by (m/p)
=p X (m/p)
=m
2) (mnx^2/y^3) divided by(x^4mn/y^3)
= (mnx^2/y^3) X (y^3)/(x^4mn)
= (mnx^2y^3)/(x^4mny^3)
=(1/x^2) (canceling mny^3)